OpenStudy (anonymous):

11* 3 sqrt 6+8* 4 sqrt 6 +2* 3sqrt 6-3* 4sqrt 6

OpenStudy (hba):

\[\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\]

OpenStudy (anonymous):

hi

OpenStudy (hba):

\[11*3\sqrt{6}+8*4\sqrt{6}+2*3\sqrt{6}-3*4\sqrt{6},Right ?\]

OpenStudy (anonymous):

A) 18 sqrt 12 B) 18 sqrt 6 C) 5 4 sqrt 12+13 3 sqrt 12 D

OpenStudy (anonymous):

D) 5 4 sqrt 6+133 sqrt 6

OpenStudy (hba):

Can you simplify it ?

OpenStudy (anonymous):

im not sure what im doing with this problem it is asking me to simplify

OpenStudy (hba):

We just have to simply add them up but before we do this we have to multiply the terms Example 2*3sqrt{6}=6sqrt{6}

OpenStudy (anonymous):

when i simplified it i got 59 sqrt 6 but thats not one of the options

OpenStudy (hba):

Show me you working :)

OpenStudy (anonymous):

let me try it your way so 11 3 sqrt 6 is ....33 sqrt 6 i think

OpenStudy (hba):

Yeah good :)

OpenStudy (hba):

Now do the others and then add,subtract them :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

33 sqrt 6+32 sqrt 6+6 sqrt 6-12 sqrt 6

OpenStudy (anonymous):

it still comes out to 59 sqrt 6

OpenStudy (hba):

Is your question correct my friend ?

OpenStudy (hba):

lol :P

OpenStudy (anonymous):

ya i think so its just im taking this class on flvs and they always have the wierdest answer ... like you solve the problem and then its wrong and apparently their answer is right

OpenStudy (hba):

Haha lol :P

OpenStudy (anonymous):

hold on the question itself is a picture ill attach it

OpenStudy (hba):

lol,the question is diffrent lemme show you this.

OpenStudy (hba):

You will add these \[\huge\ a \sqrt[3]{6}+b \sqrt[3]{6}=(a+b)\sqrt[3]{6}\] You add this and also add the other.

OpenStudy (anonymous):

i dont understand that

OpenStudy (hba):

@starrlight \[\huge\ 11\sqrt[3]{6}+8\sqrt[4]{6}+2\sqrt[3]{6}-3\sqrt[4]{6}\] Add them straight up :)

OpenStudy (hba):

@starrlight Um,getting it ?

OpenStudy (anonymous):

no

OpenStudy (hba):

\[\huge\ 11\sqrt[3]{6}+2\sqrt[3]{6}+8\sqrt[4]{6}-3\sqrt[4]{6}\]

OpenStudy (anonymous):

oh ok